a recatngular ro0m is 7ft longer than its wide.its area is 228ft squared.what is the width of the ro0m??
Answer by nerdybill(7384)

a recatngular ro0m is 7ft longer than its wide.its area is 228ft squared.what is the width of the ro0m?
.
Let w = width
then
w+7 = length
.
Area of a rectangle = length * width
.
228 = (w+7)w
228 = w^2 + 7w
0 = w^2 + 7w - 228
.
Factoring, we get:
0 = (w-19)(w+12)
w = {19, -12}
.
Since width can't be negative
w = 19 feet (width)
w+7 = 26 feet (length)